consumer surplus

A friend of your is considering two cell phone service providers. Provider A charges $120 per month for the service regadless of phone calls made. Provider B does nothave a fixed service fee but instead charges $1 per minute for calls. Your friend's monthly demand for minutes of calling is given by the equation Qd=150-50P, where Pis the price of a minute.
a. With each provider, what is the cost to your friend of an extra minute on the phone?
b. In light of your answer to (a), how many minutes would your friend talk on the phone with each provider?
c. How much would he end up paying each provider every month?
d. How much consumer surplus would be obtain with provider? (Hint: Graph the demand curve and recall the formula for the area of a triangle.)
e. Which provider would you recommend that your friend choice? Why?
Answer #1

A friend of your is considering two cell phone service providers. Provider A charges $120 per month for the service regadless of phone calls made.Provider B does not have a fixed service fee but instead charges $1 per minute for calls. Your friend's monthly demand for minutes of calling is givenby the equation Qd=150-50P, where P is the price of a minute.
a. With each provider, what is the cost to your friend of an extra minute on the phone?

Extra one minute –net work A ---------ZERO

Extra one minute –net work B ---------ONE DOLLAR

b. In light of your answer to (a), how many minutes would your friend talk on the phone with each provider?

Q=150-50P

net work A

Q=150-50P

P=0 SO Q= 150 MNUTES IN --------- net work A

Q=150-50P

net work B

Q=150-50P

P=1 SO Q= 100 MNUTES IN --------- net work A

c. How much would he end up paying each provider every month?

net work A------------- 120 DOLLARS/MONTH

net work B------------- 100 DOLLARS/MONTH


d. How much consumer surplus would be obtain with provider? (Hint: Graph the demand curve and recall the formula for the area of a triangle.)

net work A------------- ZERO surplus

net work B------------- 150 DOLLARS



answered by: ahunt
Answer #2

A = 105 B = 100